Hello Tutor,
I was hoping you could help me with this problem.
Solve the following quadratic equation by completeing the square:
2x^2 + 10x + 11 = 0
Any help this evening on this problem would be appreciated. Thank you.

Answer by stanbon(75887)

Solve the following quadratic equation by completing the square:
2x^2 + 10x + 11 = 0
2x^2 + 10x = -11
2(x^2+5x+(5/2)^2) = -11 + 2(5/2)^2
2(x+(5/2))^2 = 6/4
(x+(5/2))^2 = 3/4
x+(5/2) = +- [sqrt3]/2
x = [-5+sqrt3]/2 or x = [-5-sqrt3]/2
==================
Cheers,
Stan H.